OpenStudy (anonymous):

Given the linear equation y = 3x + 1, find the y-coordinates of the points (1, ), (-2, ), (3, ). Please show all of your work. Plot these points and graph the linear equation. Submit your graph through the dropbox

OpenStudy (chriss):

3(1)+1=4 (1,4) 3(-2)+1=-5 (-2,-5) 3(3)+1=10 (3,10)

OpenStudy (anonymous):

For the first one, substitute 1 for x and we get y = 3(1) + 1 = 4 Can you do the same for -2 and 3?